OpenStudy (anonymous):

Can some one please help me? Write an equation of the line that is parallel to the given line y = x + 3 and the given point (5,0).

OpenStudy (anonymous):

Whats the slope of the given line?

OpenStudy (anonymous):

(5,0) is x-intercept @MrDoe

OpenStudy (anonymous):

Sigh, first notice that for a parallel line the slope is the SAME. This means the slop of the line we want is just 1. That means we need something of the form: \[y-y_0=m(x-x_0)\] Where m = 1. And we have (x_0,y_0)=(5,0) \[y-0=(1)(x-5) \implies y=x-5\] The way to visualize it is to look at this drawing:  There are infinitely many lines parallel to y=x+3, we only need to use the point to find the one that fits our conditions.

OpenStudy (mrdoe):

indeed caleb, thanks for catching that, the line passing through (5,0) with a slope of one will be y=x-5

OpenStudy (anonymous):

so the slope would be .....1?

OpenStudy (anonymous):

sorry i didnt reply faster

OpenStudy (anonymous):

When you define a line in slope intercept form you have: \[y=mx+b\] Where m is the slope and b is the y-intercept (hence "slope-intercept"). So if there is a 1 or no coefficient (implied 1) in front of the x then the slope is 1.